The X-ray coming from a Coolidge tube has a cutoff wavelength of 80 pm. Find the kinetic energy of the electrons hitting the target.

Grade:11

1 Answers

Kevin Nash
9 years ago
λ = 80 pm, e = hc/ λ = 1242/80*10^3 = 15.525 * 10^3 eV = 15.5 KeV